OpenStudy (anonymous):

A brownie recipe requires four and two thirds cups of sugar to one cup of chocolate chips. If two and one third cups of sugar is used, what quantity of chocolate chips will be needed, according to the recipe?

OpenStudy (anonymous):

@ganeshie8 @Mertsj @nincompoop

OpenStudy (ivettef365):

cross multiply, what do you get

OpenStudy (anonymous):

so x=

OpenStudy (ivettef365):

do you know how to cross multiply ??

OpenStudy (anonymous):

yes i got 1/2

OpenStudy (ivettef365):

ok, lets do it together, you have 14 7 ----- x = ------- 3 3

OpenStudy (ivettef365):

that is after it is multiplied, now you have to isolate x, so you need to multiply both sides by 3/14

OpenStudy (ivettef365):

what do you get

OpenStudy (texaschic101):

4 2/3 = 14/3 2 1/3 = 7/3 (14/3)/1 = (7/3)/x -- 14/3 to 1 cup = 7/3 to x cups) cross multiply (14/3)(x) = (7/3)(1) 14/3x = 7/3 (now divide both sides by 14/3 to get x by itself) x = (7/3)/(14/3) x = 7/3 * 3/14 x = 21/42 = 1/2 lets check.. (14/3)/1 = (7/3)/(1/2) cross multiply 14/3 = (7/3)(2/1) 14/3 = 14/3 (correct) ANSWER : If 2 1/3 cups of sugar is used, then 1/2 cups of chocolate chips are used.